### v1

@@ Appendix A Detailed Notes on Individual Clusters
(26) Pal 12: This cluster is a known satellite of the Sagittarius dwarf galaxy (2010ApJ...718.1128L; 2019A&A...630L...4M). In the study by 2022A&A...665A...8K, 77 extra-tidal member stars were identified in this cluster, all of which lie beyond its Jacobi radius. However, our matched-filter S/N map shows no evidence of tail-like features associated with it. In addition, 2018MNRAS.473.3062M found no extended features based on deep photometry. It is worth noting that its tidal radius appear somewhat peculiar, with $r_{\rm t}^{\mathrm{King}}$ being much larger than $r_{\rm t}^{\mathrm{dyn}}$, and there seems to be no detailed discussion of this discrepancy in the existing literature. Therefore, we classify Pal 12 as a G3 cluster solely based on our matched-filter result, and this source is not included in the subsequent statistical analyses.
@@
(27) Pal 13: 2020AJ....160..244S identified extended tidal tails associated with this cluster using DECaLS data, and 2020A&A...635A..93P also found that its member stars extend beyond the Jacobi radius. Our matched-filter S/N map clearly reveals the presence of tidal tail structures, which are aligned with the extended features reported by 2020AJ....160..244S, although their spatial extent is less pronounced in our data. Based on these results, we classify Pal 13 as a G1 cluster.
@@
(28) NGC 7492: Our matched-filter S/N map reveals clear tidal tail structures, consistent with the findings of 2017ApJ...841L..23N based on Pan-STARRS PS1 data, and with the results of 2025AJ....170..294C. Although some studies using deeper photometric data (2018ApJ...860...66M; 2018ApJ...860...65M; 2022MNRAS.513.3136Z) did not reproduce these features, we classify NGC 7492 as a G1 cluster based on our results and the aforementioned evidence of 2017ApJ...841L..23N and 2025AJ....170..294C.

### v2

@@ Appendix A Detailed notes on individual clusters
(26) Pal 12: This cluster is a known satellite of the Sagittarius dwarf galaxy (Law and Majewski 2010; Massari et al. 2019). In the study by Kundu et al. (2022), 77 extratidal member stars were identified in this cluster, all of which lie beyond its Jacobi radius. However, our matched-filter S/N map shows no evidence of tail-like features associated with it. In addition, Musella et al. (2018) found no extended features based on deep photometry. It is worth noting that its tidal radius appear somewhat peculiar, with $r_{\rm t}^{\mathrm{King}}$ being much larger than $r_{\rm t}^{\mathrm{dyn}}$, and there seems to be no detailed discussion of this discrepancy in the existing literature. Therefore, we classify Pal 12 as a G3 cluster solely based on our matched-filter result, and this source is not included in the subsequent statistical analyses.
@@
(27) Pal 13: Shipp et al. (2020) identified extended tidal tails associated with this cluster using DECaLS data, and Piatti and Fernández-Trincado (2020) also found that its member stars extend beyond the Jacobi radius. Our matched-filter S/N map clearly reveals the presence of tidal tail structures, which are aligned with the extended features reported by Shipp et al. (2020), although their spatial extent is less pronounced in our data. Based on these results, we classify Pal 13 as a G1 cluster.
@@
(28) NGC 7492: Our matched-filter S/N map reveals clear tidal tail structures, consistent with the findings of Navarrete et al. (2017) based on Pan-STARRS PS1 data, and with the results of Chiti et al. (2025). Although some studies using deeper photometric data (Muñoz et al. 2018b, a; Zhang et al. 2022) did not reproduce these features, we classify NGC 7492 as a G1 cluster based on our results and the aforementioned evidence of Navarrete et al. (2017) and Chiti et al. (2025).
Appendix B Additional figures
Figure 7: Color-magnitude diagrams for the remaining clusters in the sample, complementing those shown in Fig. 1
Figure 8: Continued.
Figure 9: Signal-to-noise ratio distributions for the remaining clusters in the sample, complementing those shown in Fig. 2. Among them, a prominent overdensity appears below NGC 362, which is caused by contamination from the Small Magellanic Cloud. In addition, NGC 5024 and NGC 5053 have very similar CMDs and are commonly regarded as an interacting pair of clusters (Chun et al. 2010), and thus they appear in each other’s field.
Figure 10: Continued.